OpenStudy (anonymous):

A small company has 20 employees. Six of these employees will be selected randomly to be interviewed as part of an employee satisfaction program. How many different groups of six can be selected?

OpenStudy (turingtest):

I wanna see how to do this so I'm posting...

OpenStudy (anonymous):

\[\dbinom{20}{6}=\frac{20\times 19\times 18\times 17\times 16\times 15}{6\times 5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

this is a whole number so cancel first, multiply second and the calculation will not be very hard

OpenStudy (anonymous):

for example \[6\times 3=18,5\times 4=20\] so you can multiply \[19\times 17\times 8\times 15\]